# uncompress

### v2

@@ Abstract.
In this paper we investigate the degrees of irrationality of degenerations of $\epsilon$-lc Fano varieties of arbitrary dimensions. We show that given a generically $\epsilon$-lc klt Fano fibration $X\to Z$ of dimension $d$ over a smooth curve $Z$ such that $(X,tF)$ is lc for a positive real number $t$ where $F$ is the reduction of an irreducible central fibre of $X$ over a closed point $z\in Z$, then $F$ admits a rational dominant map $\pi\colon F\dashrightarrow C$ to a smooth projective variety $C$ with bounded degree of irrationality depending only on $\epsilon,d,t$ such that the general fibres of $\pi$ are irreducible and rational. This proves the generically bounded case of a conjecture proposed by the first author and Loginov for log Fano fibrations of dimensions greater than three. One of the key ingredients in our proof is to modify the generically $\epsilon$-lc klt Fano fibration $X\to Z$ to a toroidal morphism of toroidal embeddings with bounded general fibres.
@@ 1. Introduction
Bounded irrationality of divisors on Fano fibrations. Given a family of varieties parameterised by a smooth curve, degenerations are the limit varieties of this family of varieties. Many geometric properties of varieties are preserved by passing to degenerations. For example, irreducible degenerations of uniruled varieties and rationally chain connected varieties are also uniruled and rationally chain connected; see [Kol96, Theorem IV.1.8, Theorem IV.3.5.3]. However, it is well-known that rational varieties can degenerate to non-rational varieties even for rational del Pezzo surfaces; for instance, a smooth cubic surface can degenerate to a cone over an elliptic curve which is a non-rational and singular del Pezzo surface. This phenomenon is closely related to the fact that the degeneration has log canonical (lc) but not Kawamata log terminal (klt) singularities.
In the paper [BL21], the first named author and Loginov study the non-rationality property of degenerations of klt del Pezzo surfaces and prove how far the components of the degenerations can be from being rational. More precisely, the boundedness of irrationality of degenerations is proved in [BL21] as follows. In particular, by [BL21, Theorem 1.1], the irrationality of degenerations of klt del Pezzo surfaces are bounded in a certain way even though the degenerations do not belong to a bounded family of surfaces, that is, the collection of degenerations can not be parameterised by a finite type scheme; see §2.9. Recall that a contraction $f\colon X\to Z$ of normal varieties is called a klt Fano fibration if $X$ has klt singularities and $K_{X}$ is ample over $Z$; see §2.5.
Theorem 1.1 ([BL21, Theorem 1.1]).
@@
Note that the genus $g(C)$ can be arbitrarily large if $t\leq\frac{1}{2}$; see [BL21, Example 2.3]. Moreover, Theorem 1.1 follows from a more general result for Fano type log Calabi-Yau fibrations where the base $Z$ of the fibration can have dimension $\geq 1$; see [BL21, Theorem 1.3]. Recall that the gonality $\operatorname{gon}(C)$ of a smooth curve $C$ is the smallest degree of the dominant morphisms of curves $C\to\mathbb{P}^{1}$. In higher dimensions, the degree of irrationality $\operatorname{irr}(D)$ of a variety $D$ is the least possible degree of dominant rational maps $D\dashrightarrow\mathbb{P}^{\dim D}$; see §2.16. The first named author and Loginov ask in [BL21] whether a similar result to Theorem 1.1 for klt Fano fibrations of higher dimensions should also hold.
Question 1.2 (cf. [BL21, Question 1.4]).
Fix a positive real number $t>0$ and natural number $d$. Suppose that $f\colon X\to Z$ is a klt Fano fibration over a smooth curve $Z$, where $\dim X=d$. Assume that $D$ is the reduction of an irreducible fibre of $f$ such that $(X,tD)$ is lc. Is it true that there is a rational map $D\dashrightarrow C$, where the general fibres are rationally connected and $C$ is a smooth projective variety with bounded degree of irrationality?
In this paper, we apply a very different approach from [BL21] (cf. [BL21, §5, §6]) via toroidal geometry to prove the boundedness of degrees of irrationality for degenerations of klt Fano fibrations of arbitrary dimensions with $\epsilon$-lc general fibres.
@@
    $F$ is the reduction of an irreducible fibre of $f$ and $(X,tF)$ is lc.
Then there is a dominant rational map $F\dashrightarrow C$ whose general fibres are irreducible and rational, and $C$ is a bounded smooth projective variety hence with bounded degree of irrationality.
Note that the general fibres of $F\dashrightarrow C$ in Theorem 1.3 are rational, which is a much stronger geometric property than the rational connectivity predicted in Question 1.2. Example 2.1 of [BL21] shows that Theorem 1.3 does not hold without assuming that $(X,tF)$ is lc. A similar result can be formulated when the fibre containing $F$ is not irreducible by using Fano type log Calabi-Yau fibrations; cf. [BL21, Theorem 1.3, Question 1.4].
As a corollary of Theorem 1.3, we show that every irreducible component of fibres of $\epsilon$-lc Fano fibrations admits the structure as in Theorem 1.3. Note that in this case, we do not prescribe a positive real number $t>0$ so that $(X,tF)$ is lc.
@@
Let $\epsilon$ be a positive real number, and let $d$ be a natural number. Let $f\colon X\to Z$ be a Fano fibration with $\dim X=d$, where $Z$ is a smooth curve, and $X$ is $\epsilon$-lc. Let $F$ be an irreducible component of a closed fibre of $f$. Then there is a dominant rational map $F\dashrightarrow C$ whose general fibres are irreducible and rational, and $C$ is a bounded smooth projective variety hence with bounded degree of irrationality.
Relative toroidalisation and sketch of the proof. One of the key ingredients in our proof of Theorem 1.3 is that we can modify the family $f\colon X\to Z$ to a generically bounded toroidal morphism of strict toroidal embeddings. A general fibre of $X\to Z$ in Theorem 1.3 is an $\epsilon$-lc Fano variety which is bounded by [Bir21, Theorem 1.1]. Thus, there is a birational map $\phi\colon X\dashrightarrow Y$ over $Z$ such that every fibre of $g\colon Y\to Z$ is bounded, that is, $Y\to Z$ is relatively bounded; see §2.13. Moreover, $\phi$ can be chosen so that it does not contract any curve over the generic point of $Z$ as general fibres of $f$ are already bounded. In §4, we show that there exists a toroidal morphism $f^{\prime}$ fitting into the commutative diagram
@@
Note that the toroidal modification exists for any projective surjective morphism of varieties by [AK00, ADK13]; see Theorem 3.10. However, here we require a stronger condition on the boundedness of general fibres of $f^{\prime}$, which will be deduced from Theorem 3.12. Similar constructions are applied in [Bir23a] where $m_{X}$ may not be birational. The techniques related to toroidal embeddings are also used in [Jia23]; see, for example, [Jia23, §3].
Denote by $C^{\prime}$ the centre of $F$ on $Y^{\prime}$. Decreasing $t$ if necessary, we can assume that $t$ is a positive rational number. By taking an $n$-complement of $K_{X}+tF$ for some $n\in\mathbb{N}$ depending only on $t$ (see §2.6), it will be shown in §4 that the toroidal modification $f^{\prime}$ can be taken so that these additional properties are also satisfied:
@@
Let $C$ be a bounded resolution of $C^{\prime}$, then there is a rational map $\pi\colon F\dashrightarrow C$. Now, $C$ is a smooth projective variety with bounded degree of irrationality since it is bounded; cf. Lemma 2.17. Moreover, as $C^{\prime}$ is an lc centre of the strict toroidal couple $(Y^{\prime},D^{\prime})$, a general fibre of $\pi$ is irreducible and rational; see Proposition 3.5. This concludes Theorem 1.3.
The following result compares Theorem 1.3 and [BL21, Theorem 1.1] when $\dim X=3$.
Corollary 1.5 (cf. [BL21, Theorem 1.1]).
Fix positive real numbers $\epsilon,t$. Let $f\colon X\to Z$ be a klt Fano fibration where $\dim X=3$ and $\dim Z=1$ such that $X$ is $\epsilon$-lc over the generic point of $Z$. Let $F$ be the reduction of an irreducible fibre of $f$ such that $(X,tF)$ is lc. Then $F$ is birational to $\mathbb{P}^{1}\times E$, where $E$ is a smooth projective curve with gonality $\operatorname{gon}(E)$ and genus $g(E)$ bounded depending only on $\epsilon$ and $t$.
@@
Then the dimension of $C$ gives several possibilities on the structure of $F$:
@@
Note that $F$ must be ruled as it is a degeneration of a rational surface (see [Kol96, Theorem IV.1.6]), so $F$ is birational to $\mathbb{P}^{1}\times E$ for some smooth curve $E$. In case (1), we can take $E=\mathbb{P}^{1}$. In case (2), $E$ is isomorphic to $C$ which has bounded gonality and genus. Now we show that $E$ also has bounded gonality and genus in case (3). As $F$ is birationally bounded in this case, $F$ has bounded degree of irrationality by Lemma 2.17, hence $E$ also has bounded gonality which is bounded from above by the degree of irrationality of $F$; see [Bas17, Remark 2.1]. On the other hand, the irregularity of the surface $\mathbb{P}^{1}\times E$ is equal to the genus $g(E)$. Since $\mathbb{P}^{1}\times E$ is birationally bounded, there are only finitely many possible values for the irregularity of $\mathbb{P}^{1}\times E$, so $g(E)$ is also bounded from above. ∎
@@
Boundedness of $\epsilon$-lc Fano varieties. Keeping the same notation as in Theorem 1.3 and assuming that $\dim X=3$, we know that $F$ is birational to $\mathbb{P}^{1}\times E$ for some smooth projective curve $E$. Notice that the assumptions in Theorem 1.3 for $\dim X=3$ is stronger than Theorem 1.1 by requiring additionally that $X$ is $\epsilon$-lc over the generic point of $Z$, that is, a general fibre of $X\to Z$ is an $\epsilon$-lc del Pezzo surface for the fixed $\epsilon>0$. It is well-known that $\epsilon$-lc del Pezzo surfaces are bounded for any fixed $\epsilon>0$ (cf. [Ale94]), and the family of del Pezzo surfaces is not bounded for $\epsilon=0$. This leads to the phenomenon that both the gonality and genus of $E$ are always bounded under the conditions of Theorem 1.3; see Corollary 1.5. However, dropping the boundedness condition, [BL21, Example 2.3] illustrates that the genus of the curve $E$ can be arbitrarily large. A general fibre of the klt Fano fibration constructed in [BL21, Example 2.3] is isomorphic to the weighted projective space $\mathbb{P}(1,1,n)$ with $n\geq 3$. The log discrepancy of $\mathbb{P}(1,1,n)$ is equal to $2/n$, hence the family of del Pezzo surfaces $\mathbb{P}(1,1,n)$ with $n\geq 3$ is not bounded; cf. [Bir21, Example 1.2]. The boundedness condition on general fibres of the klt Fano fibration is crucial in our proof of Theorem 1.3 in higher dimensions, that is, $\epsilon$-lc Fano varieties of dimension $d$ form a bounded family for any fixed positive real number $\epsilon>0$ and natural number $d$; see [Bir21, Theorem 1.1]. It is also clear from the discussion above that boundedness of complements (see [Bir19]) is another crucial ingredient of our proofs.
@@
Acknowledgements. We thank professor Dan Abramovich for answering our numerous questions about the results in [ATW20]. We would like to thank Dingxin Zhang and Heer Zhao for very helpful discussions about logarithmic geometry. We would also like to thank Mao Sheng for his helpful comments. The first author was supported by a grant from Tsinghua University and a grant of the National Program of Overseas High Level Talent. The results and proofs of this paper were presented by the second author in a talk at the online Tsinghua Algebraic Geometry Seminar in October 2022.
@@ 2. Preliminaries
By saying a scheme, we mean a separated scheme of finite type over the fixed algebraically closed field $\mathbb{K}$ of characteristic zero. A variety is an irreducible, reduced, and quasi-projective scheme. For a scheme $X$, we denote by $X_{\text{red}}$ the maximal reduced closed subscheme of $X$ and call it the reduction of $X$.
@@ 2.1. Contractions
A contraction is a projective morphism of schemes $f:X\to Y$ such that $_{*}\mathcal{O}_{X}=\mathcal{O}_{Y}; ff$ is not necessarily birational. In particular, $f$ has connected fibres. Moreover, if $X$ is normal, then $Y$ is also normal.
@@ 2.2. Divisors
Let $X$ be a scheme. By a divisor, we mean a Weil divisor on $X$, that is, a finite $\mathbb{Z}$-linear combination of codimension one irreducible and reduced closed subschemes of $X$. By a $\mathbb{Q}$-divisor (respectively, an $\mathbb{R}$-divisor), we mean a finite linear combination $\sum_{i}b_{i}B_{i}$ where every $B_{i}$ is a prime Weil divisor on $X$ and $b_{i}\in\mathbb{Q}$ (respectively, $b_{i}\in\mathbb{R}$). A $\mathbb{Q}$-divisor (respectively, an $\mathbb{R}$-divisor) is called $\mathbb{Q}$-Cartier (respectively, $\mathbb{R}$-Cartier) if it is a $\mathbb{Q}$-linear (respectively, an $\mathbb{R}$-linear) combination of Cartier divisors. Let $B_{1}$ and $B_{2}$ be two $\mathbb{R}$-divisors on $X$. We say that $B_{1}\sim B_{2}$ (respectively, $B_{1}\sim_{\mathbb{Q}}B_{2}$, respectively, $B_{1}\sim_{\mathbb{R}}B_{2}$) if $B_{1}-B_{2}$ is a $\mathbb{Z}$-linear (respectively, a $\mathbb{Q}$-linear, respectively, an $\mathbb{R}$-linear) combination of principal divisors.
Let $f\colon X\to Z$ be a morphism of schemes, and let $L$ and $M$ be $\mathbb{R}$-divisors on $X$. We say that $L\sim M$ over $Z$ (respectively, $L\sim_{\mathbb{Q}}M$ over $Z$, respectively, $L\sim_{\mathbb{R}}M$ over $Z$) if there is a Cartier (respectively, a $\mathbb{Q}$-Cartier, respectively, an $\mathbb{R}$-Cartier) divisor $N$ on $Z$ such that $L-M\sim f^{*}N$ (respectively, $L-M\sim_{\mathbb{Q}}f^{*}N$, respectively, $L-M\sim_{\mathbb{R}}f^{*}N$).
Let $f\colon X\to Z$ be a morphism of schemes and $D$ an $\mathbb{R}$-divisor on $X$. We say that $D$ is vertical/Z/Z if $f(\operatorname{Supp}D)$ does not contain any generic point of $Z$. If $D$ does not have any vertical/Z/Z irreducible components, we say that $D$ is horizontal/Z/Z.
For volume of a big $\mathbb{R}$-divisor, we refer the readers to [Laz04, §2.2.C] for the definition and more details.

### v3

@@ Abstract.
In this paper we investigate the degrees of irrationality of degenerations of $\epsilon$-lc Fano varieties of arbitrary dimensions. We show that given a generically $\epsilon$-lc klt Fano fibration $X\to Z$ of dimension $d$ over a smooth curve $Z$ such that $(X,tF)$ is lc for a positive real number $t$, where $F$ is the reduction of an irreducible central fibre of $X$ over a closed point $z\in Z$, then $F$ admits a rational dominant map $\pi\colon F\dashrightarrow C$ to a smooth projective variety $C$ with bounded degree of irrationality depending only on $d,\epsilon,t$ such that the general fibres of $\pi$ are irreducible and rational. This proves the generically bounded case of a conjecture proposed by the first author and Loginov for log Fano fibrations of dimensions greater than three. One of the key ingredients in our proof is to modify the generically $\epsilon$-lc klt Fano fibration $X\to Z$ to a toroidal morphism of toroidal embeddings with bounded general fibres.
@@ 1. Introduction
Bounded irrationality of divisors on Fano fibrations. Given a family of varieties parameterised by a smooth curve, degenerations are the limit varieties of this family of varieties. Many geometric properties of varieties are preserved by passing to degenerations. For example, irreducible degenerations of uniruled varieties (respectively, rationally chain connected varieties) are also uniruled (respectively, rationally chain connected); see [33, Theorem IV.1.8, Theorem IV.3.5.3]. However, it is well-known that rational varieties can degenerate to non-rational varieties even for rational del Pezzo surfaces; for instance, a smooth cubic surface can degenerate to a cone over an elliptic curve which is a non-rational and singular del Pezzo surface. This phenomenon is closely related to the fact that the degeneration has log canonical (lc) but not Kawamata log terminal (klt) singularities.
In the paper [13], the first named author and Loginov study the non-rationality property of degenerations of klt del Pezzo surfaces and prove how far the irreducible components of the degenerations can be from being rational. More precisely, the boundedness of irrationality of degenerations is proved in [13] as follows. In particular, by [13, Theorem 1.1], the irrationality of degenerations of klt del Pezzo surfaces is bounded in a certain way even though the degenerations do not belong to a bounded family of surfaces, that is, the collection of degenerations can not be parameterised by a finite type scheme; see §2.9. Recall that a contraction $f\colon X\to Z$ of normal varieties is called a Fano fibration (respectively, klt Fano fibration) if $X$ has lc singularities (respectively, klt singularities) and $K_{X}$ is ample over $Z$; see §2.5.
Theorem 1.1 ([13, Theorem 1.1]).
@@
Note that the genus $g(C)$ can be arbitrarily large if $t\leq\frac{1}{2}$; see [13, Example 2.3]. Moreover, Theorem 1.1 follows from a more general result for Fano type log Calabi-Yau fibrations, where the base $Z$ of the fibration can have dimension $\geq 1$; see [13, Theorem 1.3]. Recall that the gonality $\operatorname{gon}(C)$ of a smooth curve $C$ is the smallest degree of dominant morphisms of curves $C\to\mathbb{P}^{1}$. In higher dimensions, the degree of irrationality $\operatorname{irr}(D)$ of a variety $D$ is the least possible degree of dominant rational maps $D\dashrightarrow\mathbb{P}^{\dim D}$; see §2.16. The first named author and Loginov ask in [13] whether a similar result to Theorem 1.1 for klt Fano fibrations of higher dimensions should also hold.
Question 1.2 (cf. [13, Question 1.4]).
Fix a positive real number $t>0$ and a natural number $d$. Suppose that $f\colon X\to Z$ is a klt Fano fibration over a smooth curve $Z$, where $\dim X=d$. Assume that $D$ is the reduction of an irreducible fibre of $f$ such that $(X,tD)$ is lc. Is it true that there is a rational map $D\dashrightarrow C$, where the general fibres are rationally connected and $C$ is a smooth projective variety with bounded degree of irrationality?
In this paper, we apply a very different approach from [13] via toroidal geometry to prove the boundedness of degrees of irrationality for degenerations of klt Fano fibrations of arbitrary dimensions with $\epsilon$-lc general fibres; cf. [13, §5, §6].
@@
    $F$ is the reduction of an irreducible closed fibre of $f$ and $(X,tF)$ is lc.
Then there is a dominant rational map $F\dashrightarrow C$ whose general fibres are irreducible and rational, and $C$ is a bounded smooth projective variety depending only on $d,\epsilon,t$, hence with bounded degree of irrationality.
Note that the general fibres of $F\dashrightarrow C$ in Theorem 1.3 are rational, which is a much stronger geometric property than the rational connectivity predicted in Question 1.2. Example 2.1 of [13] shows that Theorem 1.3 does not hold without assuming that $(X,tF)$ is lc. A similar result can be formulated when the fibre containing $F$ is not irreducible by using Fano type log Calabi-Yau fibrations; cf. [13, Theorem 1.3, Question 1.4].
As a corollary of Theorem 1.3, we show that every irreducible component of fibres of $\epsilon$-lc Fano fibrations admits the structure as in Theorem 1.3. Note that in this case, we do not prescribe a positive real number $t>0$ such that $(X,tF)$ is lc.
@@
Let $\epsilon$ be a positive real number, and let $d$ be a natural number. Let $f\colon X\to Z$ be a Fano fibration with $\dim X=d$, where $Z$ is a smooth curve, and $X$ is $\epsilon$-lc. Let $F$ be an irreducible component of a closed fibre of $f$. Then there is a dominant rational map $F\dashrightarrow C$ whose general fibres are irreducible and rational, and $C$ is a bounded smooth projective variety depending only on $d,\epsilon$, hence with bounded degree of irrationality.
Relative toroidalisation and sketch of the proof. One of the key ingredients in our proof of Theorem 1.3 is that we can modify the family $f\colon X\to Z$ to a generically bounded toroidal morphism of strict toroidal embeddings. A general fibre of $X\to Z$ in Theorem 1.3 is an $\epsilon$-lc Fano variety which is bounded by [8, Theorem 1.1]. Thus, there is a birational map $\phi\colon X\dashrightarrow Y$ over $Z$ such that every fibre of $g\colon Y\to Z$ is bounded; indeed, $Y\to Z$ is relatively bounded (see §2.13). Moreover, $\phi$ can be chosen so that it does not contract any curve over the generic point of $Z$ as general fibres of $f$ are already bounded. In §4, we show that there exists a toroidal morphism $f^{\prime}$ fitting into the commutative diagram
@@
Note that the toroidal modification exists for any projective surjective morphism of varieties by [1, 2]; see Theorem 3.10. However, here we require a stronger condition on the boundedness of general fibres of $f^{\prime}$, which will be deduced from Theorem 3.12. Similar constructions are applied in [11], where $m_{X},m_{Z}$ may not be birational. The techniques related to toroidal embeddings are also used in [29].
Denote by $C^{\prime}$ the centre of $F$ on $Y^{\prime}$. By decreasing $t$ if necessary, we can assume that $t$ is a positive rational number. As will be shown in §4, by taking an $n$-complement of $K_{X}+tF$ for some $n\in\mathbb{N}$ depending only on $d,t$ (see §2.6), the toroidal modification $f^{\prime}$ can be taken so that these additional properties are also satisfied:
@@
Let $C$ be a bounded nonsingular projective variety that is birational to $C^{\prime}$. Then there is a rational map $\pi\colon F\dashrightarrow C$. Now $C$ is a smooth projective variety with bounded degree of irrationality since it is bounded; cf. Lemma 2.17. Moreover, as $C^{\prime}$ is an lc centre of the strict toroidal couple $(Y^{\prime},D^{\prime})$, a general fibre of $\pi$ is irreducible and rational; see Proposition 3.5. This concludes Theorem 1.3.
The following result compares Theorem 1.3 and [13, Theorem 1.1] when $\dim X=3$.
Corollary 1.5 (cf. [13, Theorem 1.1]).
Fix positive real numbers $\epsilon,t$. Let $f\colon X\to Z$ be a klt Fano fibration, where $\dim X=3$ and $\dim Z=1$ such that $X$ is $\epsilon$-lc over the generic point of $Z$. Let $F$ be the reduction of an irreducible closed fibre of $f$ such that $(X,tF)$ is lc. Then $F$ is birational to $\mathbb{P}^{1}\times E$, where $E$ is a smooth projective curve with gonality $\operatorname{gon}(E)$ and genus $g(E)$ bounded depending only on $\epsilon$ and $t$.
@@
Then the dimension of $C$ gives several possibilities for the structure of $F$:
@@
Note that $F$ must be ruled as it is a degeneration of rational surfaces (see [33, Theorem IV.1.6]), so $F$ is birational to $\mathbb{P}^{1}\times E$ for some smooth projective curve $E$. In case (1), we can take $E=\mathbb{P}^{1}$. In case (2), $E$ is isomorphic to $C$ which has bounded gonality and genus. Now we show that $E$ also has bounded gonality and genus in case (3). As $F$ is birationally bounded in this case, $F$ has bounded degree of irrationality by Lemma 2.17, hence $E$ also has bounded gonality which is bounded from above by the degree of irrationality of $F$; see [5, Remark 2.1]. On the other hand, the irregularity of the surface $\mathbb{P}^{1}\times E$ is equal to the genus $g(E)$; see [27, Exercise III.8.4]. Since $\mathbb{P}^{1}\times E$ is birationally bounded, there are only finitely many possible values for the irregularity of $\mathbb{P}^{1}\times E$, so $g(E)$ is also bounded from above. ∎
@@
Boundedness of $\epsilon$-lc Fano varieties. Keeping the same notation as in Theorem 1.3 and assuming that $\dim X=3$, we know that $F$ is birational to $\mathbb{P}^{1}\times E$ for some smooth projective curve $E$. Notice that the assumptions in Theorem 1.3 for $\dim X=3$ are stronger than those in Theorem 1.1 by requiring additionally that $X$ is $\epsilon$-lc over the generic point of $Z$, that is, a general fibre of $X\to Z$ is an $\epsilon$-lc del Pezzo surface for the fixed $\epsilon>0$. It is well-known that $\epsilon$-lc del Pezzo surfaces are bounded for any fixed $\epsilon>0$ (cf. [4]), and the family of del Pezzo surfaces is not bounded for $\epsilon=0$. This leads to the phenomenon that both the gonality and genus of $E$ are always bounded under the conditions of Theorem 1.3; see Corollary 1.5. However, dropping the boundedness condition, [13, Example 2.3] illustrates that the genus of the curve $E$ can be arbitrarily large. A general fibre of the klt Fano fibration constructed in [13, Example 2.3] is isomorphic to the weighted projective space $\mathbb{P}(1,1,n)$ with $n\geq 3$. The log discrepancy of $\mathbb{P}(1,1,n)$ is equal to $2/n$, hence the family of del Pezzo surfaces $\mathbb{P}(1,1,n)$ with $n\geq 3$ is not bounded; cf. [8, Example 1.2]. The boundedness condition on general fibres of the klt Fano fibration is crucial in our proof of Theorem 1.3 in higher dimensions, that is, $\epsilon$-lc Fano varieties of dimension $d$ form a bounded family for any fixed positive real number $\epsilon>0$ and natural number $d$; see [8, Theorem 1.1]. It is also evident from the discussion above that boundedness of complements is another crucial ingredient of our proofs; see [7].
@@
Acknowledgements. We thank Professor Dan Abramovich for answering our numerous questions about the results in [3]. We would like to thank Dingxin Zhang and Heer Zhao for very helpful discussions about logarithmic geometry. We would also like to thank Mao Sheng for his helpful comments. We thank Hexu Liu for reading carefully a draft version of this paper, and for his many suggestions to improve the paper. We thank the anonymous referees for many helpful comments to improve this paper. The first author was supported by a grant from Tsinghua University and a grant of the National Program of Overseas High Level Talent. The results and proofs of this paper were presented by the second author in a talk at the online Tsinghua Algebraic Geometry Seminar in October 2022.
@@ 2. Preliminaries
By a scheme, we mean a separated scheme of finite type over the fixed algebraically closed field $\mathbb{K}$ of characteristic zero. A variety is an irreducible, reduced, and quasi-projective scheme; when emphasising the quasi-projectivity, we also call it a quasi-projective variety. For a scheme $X$, we denote by $X_{\text{red}}$ the maximal reduced closed subscheme of $X$ and call it the reduction of $X$.
@@ 2.1. Contractions
A contraction is a projective morphism of schemes $f\colon X\to Y$ such that $_{*}\mathcal{O}_{X}=\mathcal{O}_{Y}; ff$ is not necessarily birational. In particular, $f$ has connected fibres. Moreover, if $X$ is normal, then $Y$ is also normal.
@@ 2.2. Divisors
Let $X$ be a scheme. By a divisor, we mean a Weil divisor on $X$, that is, a finite $\mathbb{Z}$-linear combination of codimension one irreducible and reduced closed subschemes of $X$. By a $\mathbb{Q}$-divisor (respectively, an $\mathbb{R}$-divisor), we mean a finite linear combination $\sum_{i}b_{i}B_{i}$, where every $B_{i}$ is a prime Weil divisor on $X$ and $b_{i}\in\mathbb{Q}$ (respectively, $b_{i}\in\mathbb{R}$). A $\mathbb{Q}$-divisor (respectively, an $\mathbb{R}$-divisor) is called $\mathbb{Q}$-Cartier (respectively, $\mathbb{R}$-Cartier) if it is a $\mathbb{Q}$-linear (respectively, an $\mathbb{R}$-linear) combination of Cartier divisors. Let $B_{1}$ and $B_{2}$ be two $\mathbb{R}$-divisors on $X$. We say that $B_{1}\sim B_{2}$ (respectively, $B_{1}\sim_{\mathbb{Q}}B_{2}$, respectively, $B_{1}\sim_{\mathbb{R}}B_{2}$) if $B_{1}-B_{2}$ is a $\mathbb{Z}$-linear (respectively, a $\mathbb{Q}$-linear, respectively, an $\mathbb{R}$-linear) combination of principal divisors.
Let $f\colon X\to Z$ be a morphism of schemes, and let $L$ and $M$ be $\mathbb{R}$-divisors on $X$. We say that $L\sim M$ over $Z$ (respectively, $L\sim_{\mathbb{Q}}M$ over $Z$, respectively, $L\sim_{\mathbb{R}}M$ over $Z$) if there is a Cartier (respectively, a $\mathbb{Q}$-Cartier, respectively, an $\mathbb{R}$-Cartier) divisor $N$ on $Z$ such that $L-M\sim f^{*}N$ (respectively, $L-M\sim_{\mathbb{Q}}f^{*}N$, respectively, $L-M\sim_{\mathbb{R}}f^{*}N$); see [7, §2.3]. In this case, we usually write $L\sim M/Z$ (respectively, $L\sim_{\mathbb{Q}}M/Z$, respectively, $L\sim_{\mathbb{R}}M/Z$).
Let $f\colon X\to Z$ be a morphism of schemes and $D$ a nonzero $\mathbb{R}$-divisor on $X$. We say that $D$ is vertical/Z/Z if $f(\operatorname{Supp}D)$ does not contain any generic point of $Z$. If $D$ does not have any vertical/Z/Z irreducible components, we say that $D$ is horizontal/Z/Z.
For the volume of a big $\mathbb{R}$-divisor, we refer the readers to [37, §2.2.C] for the definition and more details.